# SuperCarver: Texture-Consistent 3D Geometry Super-Resolution for High-Fidelity Surface Detail Generation

### v2

@@ Abstract
Index Terms:
3D geometry super-resolution, diffusion model, mesh optimization, inverse rendering.
@@ I Introduction
Functionally, previous 3D generation frameworks are particularly designed for creating new 3D assets according to low-dimensional textual (1D-to-3D) or visual (2D-to-3D) prompts. However, there still exists a large amount of practical demands to upgrade the geometric fidelity of existing 3D assets while strictly preserving the original structural characteristics, which is a different mesh-to-mesh (3D-to-3D) task scenario.
Additionally, there also exist several lines of relevant studies, such as point cloud upsampling [18, 19], reference-based [20] and text-guided [21] 3D editing, geometric (depth/normal) map [22, 23, 24] and PBR material [25] estimation, but there is still an obvious lack of specialized explorations for 3D surface detail generation.
@@
In this paper, we make the first attempt to introduce SuperCarver, a two-stage processing framework for geometry super-resolution operating on given 3D textured meshes to generate realistic and expressive surface details that are faithful to the corresponding texture appearances, as shown in Figure 1.
@@
    We introduce the first mesh-to-mesh 3D geometry super-resolution learning framework. Our targeted task differs from previous 1D/2D-to-3D generative paradigms, offering new capabilities of upgrading the geometric quality of given mesh models in a non-destructive and identity-preserving manner.
@@ II Related Work
3D content generation [28, 29, 30] has always been a highly valuable yet rather challenging task. Unfortunately, the limited data volume and the diversity of classic 3D shape repositories [31] largely restrict the advancement of this field. In order to bypass the reliance on large-scale 3D training data, zero-shot schemes [2, 32] are designed to constrain the text-prompted 3D generation process using the off-the-shelf language-image prior of CLIP [4]. With the advent of pre-trained 2D image diffusion models [5], score distillation [3] is investigated to inject text guidance [33] into the optimization procedure of various neural field representations [34, 35]. Subsequent researches further make continuous advancements in improving generation quality and/or speed [36, 37], enabling 3D editing [21, 38], and supplementing image conditioning capabilities [39, 40]. Nevertheless, the distillation-based generation paradigm necessitates sophisticated processing to deal with inefficiency and inter-view inconsistency.
@@ II-B Inverse Rendering and Mesh Optimization
Inverse rendering aims at reasoning the underlying physical attributes of a target 3D scene from its 2D image observations. Different 3D scene parameters such as geometry, texture, material, and lighting can be optimized via backpropagating the gradients w.r.t. the 2D rendering outputs [57, 58]. Earlier research investigates various differentiable renderers [59, 60]. Later studies [61, 62, 63] devote to learning neural field representations. For gradient-based mesh optimization, the most straightforward way is to explicitly perform surface deformation from pre-defined mesh templates [64, 65]. Nevertheless, directly optimizing the mesh vertex positions is prone to surface defect and degeneracy, thus requiring careful initialization, regularization, and remeshing [66, 67, 68]. Another choice is to employ occupancy and distance fields [69, 70, 71] for implicit surface modeling. Since traditional iso-surface extraction algorithms [72, 73] cannot propagate gradients, various differentiable schemes [74, 75] are designed to operate on discrete occupancy grids (DMC [76]), continuous signed distance fields (MeshSDF [77]), or deformable tetrahedrons (DMTet [78]). Moreover, NJF [79] depicts the mappings between pairs of source and target triangles via Jacobian matrices. FlexiCubes [80] incorporates carefully-chosen learnable parameters into [73] for achieving much more flexible local adjustment of mesh vertices.
@@ III-B Deterministic Prior-Guided Normal Diffusion
Straightforward Solution. To achieve this goal, a straightforward diffusion-based solution is to denoise the noisy latent representation of the target normal map $\mathcal{H}$ while incorporating the color image $\mathcal{I}$ into the denoising U-Net. Thus, the forward process can be formulated as:
~$\widetilde{\mathcal{H}}_{t}=\sqrt{\bar{\alpha}_{t}}\cdot\widetilde{\mathcal{H}}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon,\quad t=\{1,...,T\},$ (1)
where the noise schedule term $\bar{\alpha}_{t}$ controls the transformations to the standard Gaussian distribution $\epsilon$, $T$ is the number of time steps, $\widetilde{\mathcal{H}}$ denotes the latent representation of $\mathcal{H}$ after VAE’s encoder, and ~$\widetilde{\mathcal{H}}_{t}$ is the noisy feature map.
@@
where $\mathbf{c}$ is the vectorized embedding originally extracted from textual condition, and the visual cues can be injected by means of concatenating ~$\widetilde{\mathcal{H}}_{t}$ with $\widetilde{\mathcal{I}}$, the encoded 2D latent feature map of the color image $\mathcal{I}$.
@@
~$\widetilde{\mathcal{H}}_{t}=\sqrt{\bar{\alpha}_{t}}\cdot\widetilde{\mathcal{H}}+\sqrt{1-\bar{\alpha}_{t}}\cdot\widetilde{\mathcal{N}},\quad t=\{1,...,T\},$ (3)
@@
Thus, we can achieve prior-guided diffusion by concatenating $\widetilde{\mathcal{S}}$ with the noisy latent representation ~$\widetilde{\mathcal{H}}_{t}$. The corresponding optimization objective can be formulated as:
@@ III-C Noise-Resistant Inverse Normal Rendering
We introduce a simple yet effective distance field deformation approach, which works in a “value-fixing, grid-moving” manner. Specifically, assuming that the original mesh $\mathcal{M}=(\mathcal{V},\mathcal{F})$ has already been normalized into the unit sphere, we start by constructing uniform 3D grid points $\mathcal{G}\in\mathbb{R}^{r\times r\times r\times 3}$ within $[-1,1]^{3}$ and computing per-grid signed distance values $\mathcal{X}\in\mathbb{R}^{r\times r\times r}$. Different from common ways of implementing deformable distance field representations via only updating $\mathcal{X}$ or jointly updating both $\mathcal{G}$ and $\mathcal{X}$, we choose a subtle strategy that fixes the distance values $\mathcal{X}$ while separately learning zero-initialized offsets $\mathcal{O}\in\mathbb{R}^{r\times r\times r\times 3}$ to move initial grid positions $\mathcal{G}$. In comparison, standard distance field refinement requires higher resolutions of fixed uniform grids; otherwise, detailed structures cannot be well captured even with accurate distance values. Additionally, fixing pre-computed distance values hinders the optimization process from drastically changing the original mesh structure even with obviously erroneous normal prediction in certain areas.
@@ IV Experiments
This section begins by introducing our main technical implementations in terms of data preparation and framework hyperparameters. For performance evaluations, we qualitatively illustrate the boosting effects of applying the proposed SuperCarver to a collection of diverse real-world 3D assets as well as synthesized mesh models from different 3D AIGC methods. To validate the superiority of our core technical components, we make comparisons with state-of-the-art normal prediction and mesh optimization approaches. We further conduct a series of ablation studies and analyses. In the end, we also discuss our current limitations and potential future explorations.
@@
Implementation Details. We employed the popular diffusion architecture of Stable Diffusion V2.1 [5] and performed fine-tuning on 16 NVIDIA Tesla V100 GPUs for 40K steps with a total batch size of 128. We applied the AdamW [91] optimizer with cosine annealing, reducing the learning rate from 5e-5 to 1e-5. For both training and inference, the image resolution is uniformly configured to 512$\times$512.
For iterative mesh optimization, we adopted Nvdiffrast [92] for efficient differentiable rendering from 12 camera poses (8 evenly-spaced horizontal views with azimuth angles from 0 to 360 and 4 downward-looking (elevated by 30) views with azimuth angles [45, 135, 225, 315]). Camera intrinsics were kept the same as our configurations in BlenderProc. The hyperparameters were empirically tuned as $r=512$, $\tau=0.5$, $\omega_{s}=0.25$, $\omega_{n}=0.01$ for 200 iterations.
When deployed on NVIDIA L20 GPU (mid-range product for AI inference), the average time consumption for the whole geometry super-resolution workflow is around 1$\sim$2 minutes, which is thought to be adequately efficient.
@@ IV-A Performance Evaluation
Figure 7: Typical surface damages produced by FlexiCubes [80].
@@
Considering the lack of specialized researches on our targeted task of texture-consistent geometry super-resolution, we provided qualitative visualizations for the outputs of applying SuperCarver to diverse geometrically-coarse textured meshes. As illustrated in Figure 3, our approach is capable of supplementing realistic and expressive surface details under a variety of testing cases.
In addition to real-world 3D assets, we further experimented with AI-generated mesh assets by applying SuperCarver to the textured mesh models generated from advanced 3D generation frameworks. We adopted InstantMesh [9], Unique3D [12], and CLAY [13], covering representative generation paradigms of feed-forward large reconstruction models, multi-view diffusion models, and native 3D latent diffusion models. As compared in Figure 4, various input meshes are effectively boosted by our approach to produce realistic and expressive 3D surface details well aligned with the corresponding texture appearances.
Comparison on Normal Prediction. To evaluate the effectiveness of our specialized normal diffusion architecture, we made comparisons with two advanced normal diffusion frameworks, including GeoWizard [24] and StableNormal [88]. To facilitate quantitative evaluations, we manually selected 50 detail-rich textured meshes as ground-truths from the widely-used Objaverse [6] and ABO [93] shape repositories. We applied surface simplification and smoothing to produce geometrically-coarse meshes as the input samples.
As compared in Figure 5, our normal map predictions show better details and avoid obviously erroneous areas. Besides, we further deduced the metric of mean angular error (MAE) between ground-truth and predicted normal maps, where we masked invalid backgrounds for the outputs of GeoWizard and StableNormal. As shown in Table I, thanks to our specialized diffusion architecture, our normal map predictions significantly outperform the competing approaches.
TABLE I: Quantitative comparison of normal map prediction accuracy under the metric of mean angular error (MAE).
@@
TABLE II: Normal map diffusion accuracy of different baseline schemes. “Stoc.” denotes the straightforward implementation of stochastic diffusion, then “w/o Sem.” and “w/o Geo.” respectively represent removing the high-level semantics and low-level depth priors.
@@
TABLE III: Quantitative comparison of our different baseline design choices for achieving distance field deformation. “Only Grids” represents the original design, “Only Dist.” and “Joint G&D” are the corresponding two variants.
@@
TABLE IV: Quantitative evaluation of cross-view consistency under the metric of mean angular error (MAE).
@@
Comparison on Mesh Optimization. To facilitate 3D surface geometry updating, many recent mesh generation frameworks [94, 95, 96, 12] introduce the CR [68] algorithm with joint optimization and remeshing, which is regarded as the most relevant competitor. In fact, we had previously attempted to adapt FlexiCubes [80], an advanced iso-surface representation, to achieve gradient-based geometry refinement starting from the initially given mesh surface. However, through extensive attempts, we empirically found that it easily leads to severe surface damages, as illustrated in Figure 7. We reason that this is attributable to its significantly increased deformation flexibility (which should have been its major strength) amplifies the impact of erroneous normal pixels. As shown in Figure 8, for typical hard cases with less accurate normal predictions, CR tends to produce noisy and degraded surfaces.

### v3

@@ I Introduction
Functionally, previous 3D generation frameworks are particularly designed for creating new 3D assets according to low-dimensional textual (1D-to-3D) or visual (2D-to-3D) prompts. However, there still exists a large amount of practical demands to upgrade the geometric fidelity of existing 3D assets while strictly preserving the original structural characteristics, which is a different mesh-to-mesh (3D-to-3D) task scenario. Besides, there exist several lines of relevant studies, such as point cloud upsampling [18, 19], reference-based [20] or text-guided [21] 3D editing, dense geometric map (i.e., depth/normal) [22, 23, 24] or PBR material [25] estimation, but none of them are suitable for our targeted surface detail generation. More recent works [26] devote to 3D geometry enhancement, but the detailization performance and the identity-preserving capability are still far from satisfactory.
@@
This paper introduces SuperCarver, a two-stage processing framework for geometry super-resolution operating on given 3D textured meshes to generate realistic and expressive surface details that are faithful to the corresponding texture appearances, as shown in Figure 1.
@@
    We introduce a novel mesh-to-mesh 3D geometry super-resolution learning framework. Our targeted task differs from previous 1D/2D-to-3D generative paradigms, offering new capabilities of upgrading the geometric quality of given mesh models in a non-destructive and identity-preserving manner.
@@ II Related Work
3D content generation [29, 30, 31, 32, 33, 34] has always been a highly valuable yet rather challenging task. Unfortunately, the limited data volume and the diversity of classic 3D shape repositories [35] largely restrict the advancement of this field. In order to bypass the reliance on large-scale 3D training data, zero-shot schemes [2, 36] are designed to constrain the text-prompted 3D generation process using the off-the-shelf language-image prior of CLIP [4]. With the advent of pre-trained 2D diffusion models [5], score distillation [3] is investigated to inject text guidance [37, 38, 39] into the optimization procedure of various neural field representations [40, 41]. Subsequent researches further make continuous advancements in improving generation quality and/or speed [42, 43], enabling 3D editing [21, 44], and supplementing image conditioning capabilities [45, 46]. Nevertheless, the distillation-based generation paradigm necessitates sophisticated processing to deal with inefficiency and inter-view inconsistency.
@@ II-B Inverse Rendering and Mesh Optimization
Inverse rendering aims at reasoning the underlying physical attributes of a target 3D scene from its 2D image observations. Different 3D scene parameters, including geometry, textures, materials, and lighting, can be optimized through backpropagating the gradients w.r.t. the 2D rendering outputs [63, 64]. Earlier research investigates various differentiable renderers [65, 66]. Later studies [67, 68, 69] devote to learning neural field representations. For gradient-based mesh optimization, the most straightforward strategy is to explicitly perform surface deformation from pre-defined mesh templates [70, 71]. Nevertheless, directly optimizing the mesh vertex positions can be prone to surface defect and degeneracy, thus requiring careful initialization, regularization, and remeshing [72, 73, 74]. Another choice is to employ occupancy and distance fields [75, 76, 77] for implicit surface modeling. Since traditional iso-surface extraction algorithms [78, 79] cannot propagate gradients, various differentiable schemes [80, 81] are designed to operate on discrete occupancy grids (DMC [82]), continuous signed distance fields (MeshSDF [83]), and deformable tetrahedrons (DMTet [84]). Moreover, NJF [85] depicts the mappings between pairs of source and target triangles through Jacobian matrices. FlexiCubes [86] incorporates the carefully-chosen learnable parameters into [79] for achieving much more flexible local adjustment of mesh vertices.
@@
Moreover, Sherpa3D [91] exploits a coarse 3D prior to guide a 2D diffusion model for mesh refinement. CraftsMan3D [92] incorporates a normal-driven mesh geometry refinement stage, in which new surface details can be hallucinated from coarse normal maps. DetailGen3D [26] achieves geometry enhancement through training a rectified flow [93] model that maps the latent codes between coarse and fine geometries. Despite the current progress, it is still challenging to generate high-fidelity surface details and achieve geometry-texture consistency while effectively preserving the original structure.
@@ III-B Deterministic Prior-Guided Normal Diffusion
Diffusion probabilistic models [94] define a forward Markov chain gradually perturbing data into noise, and learn a reverse process reconstructing data from Gaussian noise via iterative denoising steps.
@@
Straightforward Solution. To achieve this goal, one straightforward diffusion-based solution is to denoise the noisy latent representation of the target normal map $\mathcal{H}$ while incorporating the color image $\mathcal{I}$ into the denoising U-Net. Thus, the forward process can be formulated as:
$\widetilde{\mathcal{H}}_{t}=\sqrt{\bar{\alpha}_{t}}\cdot\widetilde{\mathcal{H}}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon,\quad t=\{1,...,T\},$ (1)
where the noise schedule term $\bar{\alpha}_{t}$ controls the transformations to the standard Gaussian distribution $\epsilon$, $T$ is the number of time steps, $\widetilde{\mathcal{H}}$ denotes the latent representation of $\mathcal{H}$ after VAE’s encoder, and $\widetilde{\mathcal{H}}_{t}$ is the noisy feature map.
@@
where $\mathbf{c}$ is the vectorized embedding originally extracted from textual condition, and the visual cues can be injected by means of concatenating $\widetilde{\mathcal{H}}_{t}$ with $\widetilde{\mathcal{I}}$, the encoded 2D latent feature map of the color image $\mathcal{I}$.
@@
$\widetilde{\mathcal{H}}_{t}=\sqrt{\bar{\alpha}_{t}}\cdot\widetilde{\mathcal{H}}+\sqrt{1-\bar{\alpha}_{t}}\cdot\widetilde{\mathcal{N}},\quad t=\{1,...,T\},$ (3)
@@
Thus, we can achieve prior-guided diffusion by concatenating $\widetilde{\mathcal{S}}$ with the noisy latent representation $\widetilde{\mathcal{H}}_{t}$. The corresponding optimization objective can be formulated as:
@@ III-C Noise-Resistant Inverse Normal Rendering
Figure 5: Results of applying our proposed SuperCarver on textured CAD mesh models collected from ShapeNet [35].
By mathematically inverting the process of image formation, inverse rendering aims at estimating the underlying physical attributes of the target 3D scene, including geometry, textures, materials, and lighting, from 2D image observations.
@@
Figure 6: Comparison of geometry enhancement results produced by (b) DetailGen3D [26] and (c) SuperCarver. The left column presents the (a) input textured meshes.
We introduce a simple yet effective distance field deformation approach, which works in a “value-fixing, grid-moving” manner. Specifically, assuming that the original mesh $\mathcal{M}=(\mathcal{V},\mathcal{F})$ has already been normalized into the unit sphere, we start by constructing uniform 3D grid points $\mathcal{G}\in\mathbb{R}^{r\times r\times r\times 3}$ within $1,1]^{3}$ and computing per-grid signed distance values $\mathcal{X}\in\mathbb{R}^{r\times r\times r}$. Different from common ways of implementing deformable distance field representations via only updating $\mathcal{X}$ or jointly updating both $\mathcal{G}$ and $\mathcal{X}$, we choose a subtle strategy that fixes the distance values $\mathcal{X}$ while separately learning zero-initialized offsets $\mathcal{O}\in\mathbb{R}^{r\times r\times r\times 3}$ to move initial grid positions $\mathcal{G}$. In comparison, standard distance field refinement requires higher resolutions of fixed uniform grids; otherwise, detailed structures cannot be well captured even with accurate distance values. Additionally, fixing pre-computed distance values hinders the optimization process from drastically changing the original mesh structure even with obviously erroneous normal prediction in certain areas.
@@
Figure 7: Comparison of geometry enhancement results produced by (b) DetailGen3D [26] and (c) SuperCarver when evaluating more diverse distributions (e.g., humanoids, non-rigid, organic, scene scan) of the (a) input textured meshes.
@@ IV Experiments
This section begins by introducing our main technical implementations in terms of data preparation and framework hyperparameters. For performance evaluations, we qualitatively illustrate the boosting effects of applying the proposed SuperCarver to a collection of diverse real-world 3D assets as well as synthesized mesh models from different data distributions. Comprehensive comparisons are conducted to show the advantages of our framework against other relevant approaches. To validate the superiority of our core technical components, we make comparisons with state-of-the-art normal prediction and mesh optimization approaches. We further conduct a series of ablation studies and analyses. In the end, we also discuss our current limitations and potential future explorations.
@@
Implementation Details. We employed the popular diffusion architecture of Stable Diffusion V2.1 [5] and performed fine-tuning on 16 NVIDIA Tesla V100 GPUs for 40K steps with a total batch size of 128. We applied the AdamW [101] optimizer with cosine annealing, reducing the learning rate from 5e-5 to 1e-5. For both training and inference, the image resolution is uniformly configured to $\times512$.
For iterative triangular mesh optimization, we adopted Nvdiffrast [102] for efficient differentiable rendering from 12 camera poses (8 evenly-spaced horizontal views with azimuth angles from 0 to 360 and 4 downward-looking (elevated by 30) views with azimuth angles [45, 135, 225, 315]). Camera intrinsics were kept the same as our configurations in BlenderProc. The hyperparameters were empirically tuned as $r=512$, $\tau=0.5$, $\omega_{s}=0.25$, $\omega_{n}=0.01$ for 200 iterations.
[…]
To facilitate quantitative evaluations, we manually selected 50 detail-rich textured meshes as our ground-truths from the widely-used Objaverse [6] and ABO [103] shape repositories. We applied surface simplification and smoothing to produce geometrically-coarse meshes as the input samples.
@@ IV-A Performance Evaluation
Figure 10: Typical surface damages produced by the optimization process of FlexiCubes [86].
@@
We presented the qualitative visualizations for the outputs of applying SuperCarver to diverse geometrically-coarse textured meshes. As illustrated in Figure 3, our approach is capable of supplementing realistic and expressive surface details under a variety of testing cases.
In addition to real-world 3D assets, we further experimented with AI-generated mesh assets by applying SuperCarver to the textured mesh models generated from advanced 3D generation frameworks. We adopted InstantMesh [9], Unique3D [12], and CLAY [13], covering representative generation paradigms of feed-forward large reconstruction models, multi-view diffusion models, and native 3D latent diffusion models. As compared in Figure 4, various input meshes are effectively boosted by our approach to produce realistic and expressive 3D surface details well aligned with texture appearances. Furthermore, we evaluated textured CAD meshes in the classic ShapeNet [35] repository (note that typical textureless CAD meshes such as those in the classic ABC [104] dataset are not the type of input our approach is designed for). As illustrated in Figure 5, our approach successfully produces high-fidelity detail patterns.
In particular, we made comprehensive comparisons with the more recent competing approach of DetailGen3D [26] targeted at 3D geometry enhancement. As depicted in Figures 6 and 7, DetailGen3D cannot sufficiently preserve the original structure of input meshes and the generated geometric patterns are much less expressive, because its image guidance effects are rather weak and ambiguous. We further conducted quantitative evaluations in Table I, where our approach consistently outperforms DetailGen3D across all the involved 2D (MAE) and 3D (CD, F-Score) metrics. On average, DetailGen3D requires around $1$ minute for geometry enhancement, while our framework takes around $1.5$ minutes. Although our framework design involves iterative optimization, its actual running efficiency is basically satisfactory and comparable to DetailGen3D.
TABLE I: Quantitative comparison of geometry enhancement results produced by DetailGen3D [26] and SuperCarver, including the 2D metric of MAE, the 3D metrics of Chamfer Distance (CD) and F-Score (following the implementation adopted in [70]), as well as the overall running efficiency (evaluated on the same device environment with an NVIDIA L20 GPU).
   Method    MAE $\downarrow$    CD ($\times 10^{-4}$) $\downarrow$    F-Score $\uparrow$    Time Cost
   DetailGen3D [26]    9.11    1.84    57.31%    $\sim\,$58s
   Ours    4.22    0.41    86.18%    $\sim\,$95s
Comparison on Normal Prediction. To evaluate the effectiveness of our specialized normal diffusion architecture, we made comparisons with two advanced normal diffusion frameworks, including GeoWizard [24] and StableNormal [100].
As compared in Figure 8, our normal map predictions show better details and avoid obviously erroneous areas. Besides, we further deduced the metric of mean angular error (MAE) between ground-truth and predicted normal maps, where we masked invalid backgrounds for the output images of GeoWizard and StableNormal. As compared in Table II, thanks to our specialized diffusion architecture, our normal map predictions significantly outperform the competing approaches.
TABLE II: Quantitative comparison of normal map prediction accuracy under the metric of mean angular error (MAE).
@@
TABLE III: Normal map diffusion accuracy of different baseline schemes. “Stoc.” denotes the straightforward implementation of stochastic diffusion, then “w/o Sem.” and “w/o Geo.” respectively represent removing the high-level semantics and low-level depth priors.
@@
TABLE IV: Quantitative comparison of our different baseline design choices for achieving distance field deformation. “Only Grids” represents the original design, “Only Dist.” and “Joint G&D” are the corresponding two variants.
@@
TABLE V: Quantitative evaluation of cross-view consistency under the metric of mean angular error (MAE).
@@
Comparison on Mesh Optimization. To facilitate 3D surface geometry updating, many recent mesh generation frameworks [105, 106, 107, 12] introduce the CR [74] algorithm with joint optimization and remeshing, which is regarded as the most relevant competitor. In fact, we had previously attempted to adapt FlexiCubes [86], an advanced iso-surface representation, to achieve gradient-based geometry refinement starting from the initially given mesh surface. However, through extensive attempts, we empirically found that it easily leads to severe surface damages, as illustrated in Figure 10. We reason that this is attributable to its significantly increased deformation flexibility (which should have been its major strength) amplifies the impacts of erroneous normal pixels. As shown in Figure 11, for typical hard cases with less accurate normal predictions, CR tends to produce noisy and degraded surfaces.
Figure 12: Illustration of prediction variation of stochastic baseline diffusion as displayed in (d). The color image, ground-truth normal map, and the output of our original design are presented in (a), (b), and (c).
Figure 13: Robustness of our distance field deformation approach when deforming the same source mesh to target meshes with increasingly larger deformation levels, with the corresponding Chamfer Distance (CD, $\times 10^{-4}$) measurement reported.